# Breaking Memory Limits: Gradient Wavelet Transform Enhances LLMs Training

### v1

@@ 4.2 Adam with GWT
Here, are the decay rates hyperparameters which are set as commonly, to preserve numerical stability, denotes the batch gradient at time step , represents the element-wise multiplication and division, and represents the first/second-order moment in Adam optimizer.
@@
For ease of presentation, we adopt a 1-level DHT in the following discussion. To integrate GWT with the Adam optimizer, we begin by applying GWT to the gradient matrix at time step (Eq. (3)), focusing on the approximate coefficients. This results in a compressed gradient matrix (approximation coefficient) and a detail coefficient matrix . We then update the Adam states by replacing the original gradient in the update rule (Eq. (5)) with the smaller matrix to calculate . To preserve the consistency of the state update, we scale the wavelet detail coefficients by dividing them by by based on the linear properties of DHT (Eq. (3)). Finally, project back to the original space and update the weights. The update rule can be rewritten by the following
@@
The procedure for applying GWT to the Adam optimizer is outlined in Algorithm 1. We implement the Adam optimizer with Haar-2 on the LLaMA 60M and 130M models, and present the learning curve in Figure 5. We observe that the learning curve (blue) demonstrates a heavy spike around the 600th iteration in LLaMA 60M, and a twice spike in LLaMA 130M. This phenomenon is likely due to unstable gradients during the early stages of training, as observed in previous studies [35, 14]. To address this issue, we adopt the Norm-growth Limiter (NL) technique introduced by Fira [14], which stabilizes gradient growth. NL controls the ratio of the current gradient norm to the previous gradient norm by rescaling the current gradient as follows
@@
where is a threshold that limits the gradient norm growth, we adopt by default. This technique is more effective than gradient clipping and significantly reduces the impact of unstable gradients in the early training stages. After introducing NL, the learning curve (orange) for our Haar-2 implementation becomes smoother, resulting in a lower training loss (from 3.41 to 3.37 in LLaMA 60M).
In summary, we integrate the GWT method into the Adam optimizer, and it is theoretically observed that the memory occupied by the optimizer state can be reduced by up to 75%. GWT leverages a memory-efficient approach by projecting gradients into a subspace and performing optimizer state updates within that subspace. This reduces memory usage and can be applied to most optimizers that typically require significant memory for storing their states. Next, we will present a more general framework to incorporate the GWT method into a wider range of optimizers.
4.3 General Optimizer with GWT
This equation captures the integration of wavelet transform-based adjustments into the optimizer’s update rule. To this end, we have developed a unified update schedule for optimizers with GWT based on linear formulation and demonstrated that GWT can be directly applied to existing optimizers with heavy-state memory usage. We present the unified framework for applying our GWT to optimizers in Algorithm 2. Therefore, we incorporate our GWT into Adafactor [27] optimizer for LLaMA pre-training tasks, and the experimental results are shown in Figure 6.
4.4 Additional Hyperparameters
In addition to the hyperparameters required by the original optimizer, several additional hyperparameters are needed to apply our GWT method. The hyperparameter stands for the scale factor. Specifically, the scale hyperparameter determines the magnitude of the update for the GWT module of ( for the modules without GWT). The hyperparameters depend on the GWT basis used in the process (e.g., Haar, DB2, Rbio2.6, etc.). The choice of filter impacts the transformation applied during the wavelet decomposition and reconstruction. stands for the GWT level, 1 denotes 1-level GWT which can reduce the optimizer memory up to 50%, and 2 denotes 2-level GWT which reduces the memory usage up to 75%, respectively. These levels allow for varying degrees of memory reduction depending on the depth of the wavelet decomposition applied. Larger can bring higher memory reduction but influence the throughput. These additional hyperparameters allow for more flexibility in applying GWT to different optimization problems while controlling the trade-off between memory usage and update precision.
@@ A.1 Network Architecture
| Model | Params | Hidden | Intermediate | Heads | Layers | Iteration | Tokens |
@@ A.2 Experiment Hyperparameters
For the GWT method, we perform hyperparameter tuning on the scale parameter from the set , selecting the value that yields the best performance across all LLaMA pre-training tasks. We found that our method is not sensitive to scale, and setting to 0.25 yields good results in most pretraining tasks, which is consistent with the hyperparameters used in previous work [12, 19]. For the baseline methods, we use their default hyperparameters: and for GaLore, and for Adam (0.0005 for Adam in LLaMA 1b). Consistent with prior work [12], we enable GWT and GaLore in the attention and MLP modules.
@@
| 60M | 130M | 350M | 1B | |
| Haar-2 | 0.25 | 0.25 | 0.25 | 0.25 |
| Haar-3 | 0.25 | 0.25 | 0.25 | - |
| DB2-2 | 0.25 | 0.25 | 0.25 | 0.25 |
| DB2-3 | 0.25 | 0.25 | 0.25 | - |
| GaLore | 0.25 | 0.25 | 0.25 | 0.25 |
| GPUs (3090) | 4 | 4 | 16 | 32 |
| Time | 0.99h | 4.03h | 11.83 | 40.18h |

### v3

@@ 3.2 Adam with GWT
Here, are the decay rates hyperparameters which are set as commonly, to preserve numerical stability, denotes the batch gradient at time step , and represents the first/second-order moment in Adam optimizer.
@@
For ease of presentation, we adopt a 1-level DHT in the following discussion. To integrate GWT with the Adam optimizer, we begin by applying GWT to the gradient matrix at time step (Eq. (2)), focusing on the approximate coefficients. This results in a compressed gradient matrix (approximation coefficient) and a detail coefficient matrix . We then update the Adam states by replacing the original gradient in the update rule (Eq. (5)) with the smaller matrix to calculate . To preserve the consistency of the state update, we scale the wavelet detail coefficients by dividing them by by based on the linear properties of DHT (Eq. (2)). Finally, project back to the original space and update the weights. The update rule can be rewritten as follows if we have ,
then compute the normalized coefficients by
Therefore, we can update the weight in the original space
@@
The procedure for applying GWT to the Adam optimizer is outlined in Algorithm 1. We implement the Adam optimizer with Haar-2 on the LLaMA 130M model, and present the learning curve in Figure 3. We observe that the learning curve (blue) demonstrates a twice spike around the 900th and 2500th iteration. This phenomenon is likely due to unstable gradients during the early stages of training, as observed in previous studies [6, 30]. To address this issue, we adopt the Norm-growth Limiter (NL) technique introduced by Fira [6], which stabilizes gradient growth. NL controls the ratio of the current gradient norm to the previous gradient norm by rescaling the current gradient as follows
@@
where is a threshold that limits the gradient norm growth, we adopt by default. This technique significantly reduces the impact of unstable gradients in the early training stages [56]. After introducing NL, the learning curve (orange) for our Haar-2 implementation becomes smoother, resulting in a lower training loss.
In summary, we integrate the proposed GWT method into the Adam optimizer and observe that it can theoretically reduce the memory footprint of optimizer states by up to 75%. GWT achieves this by projecting gradients into a lower-dimensional subspace and performing updates to the optimizer states within this subspace. This memory-efficient strategy significantly reduces the storage requirements and can be applied to a wide range of optimizers that maintain substantial state information.
Importantly, the GWT is not limited to Adam. In Appendix F, we present a unified formulation for extending GWT to other memory-intensive optimizers. We also provide empirical results demonstrating that GWT, when integrated with optimizers such as Adam-mini [53] and MUON [24], achieves comparable performance to those optimizers.
4 Experiments
In this section, we demonstrate that the Gradient Wavelet Transform (GWT), as a memory-efficient optimization technique, can achieve performance comparable to or even surpassing that of full-rank optimizers, while significantly reducing memory usage and increasing training throughput. Specifically, we evaluate GWT in two contexts: pre-training LLaMA [43] models on the Colossal Clean Crawled Corpus (C4) [34] English benchmark, fine-tuning various pre-trained models on the Multi-task Language Understanding (MMLU) [18] and RoBERTa-base on the General Language Understanding Evaluation (GLUE) [45] benchmarks. We adopt the BF-16 format to reduce memory usage across all experiments and provide a detailed description of these datasets in Appendix G.
For GWT, we use the discrete Haar wavelet as the default filter. Additionally, we present an ablation study to examine the impact of key hyperparameters, including the scale factor (), initial learning rate (), the GWT level (), and other wavelet bases, as detailed in Appendix A.2. A comprehensive description of the experimental setup can be found in Appendix A.
| Methods | 60M | 130M | 350M | 1B |
| Full-Rank Adam | 33.37 (0.36G) | 25.08 (0.76G) | 18.75 (2.06G) | 16.10 (7.80G) |
| GWT-2 | 29.35 (0.24G) | 22.47 (0.52G) | 16.29 (1.22G) | 13.50 (4.38G) |
| APOLLO-1/4 | 31.53 (0.24G) | 23.35 (0.52G) | 16.73 (1.22G) | 14.10 (4.38G) |
| GWT-3 | 29.81 (0.22G) | 22.63 (0.48G) | 16.35 (1.08G) | 13.48 (3.81G) |
| GaLore-1/8 | 48.48 (0.22G) | 30.02 (0.48G) | 21.59 (1.08G) | 17.52 (3.81G) |
| APOLLO-1/8 | 32.50 (0.22G) | 23.74 (0.48G) | 16.98 (1.08G) | 14.32 (3.81G) |
| Training Tokens | 1.3B | 2.6B | 7.8B | 13.1B |
4.1 Memory-Efficient Pre-Training
We integrate our proposed GWT method into the Adam optimizer [22], using default hyperparameters (, , ), and evaluate its performance on the LLaMA models using the C4 pretraining task. To provide a comprehensive comparison, we include our reproduced results for full-rank Adam [22], GaLore [54], and APOLLO [56], alongside reference results reported by Fira [6], and LoRA [54]. All experiments adopt a consistent configuration: we use GaLore’s default hyperparameters (, ) for both GaLore and GWT, and for APOLLO. A cosine annealing schedule is applied to the learning rate. We use a sequence length of 256 and a batch size of 512, resulting in a total of 131K tokens per batch, unless stated otherwise.
For GaLore-r and APOLLO-r, we experiment with compression ratios corresponding to 1/4 and 1/8 of the model rank, denoted as GaLore-1/4 and GaLore-1/8, respectively. These settings are aligned with GWT configurations using 2- and 3-level decompositions. Following prior work [54], we apply the same network initialization strategy and enable GWT in both the Multi-Layer Perceptron (MLP) and attention modules of the transformer.
GWT is effective in pre-training. The final validation perplexity (PPL) and estimated memory usage for all methods are summarized in Table 1. Experimental results show that GWT consistently achieves lower validation PPL while reducing memory usage compared to other memory-efficient baselines. In all the pre-training cases, GWT outperforms full-rank optimizers. Specifically, in the LLaMA 1B pre-training experiment, our 2-level GWT method reduces the total memory cost by 51% and the optimizer memory cost by 76% (from 7.80GB to 3.81GB on LLaMA 1B). Even with the use of 2-level GWT, our method still achieves experimental results that surpass full-rank, while reducing total memory consumption by up to 43%. In the same scenario, the gap between GaLore-1/8 and full-rank is further widened. Details of the memory estimation are presented in Appendix A.4.
GWT leads to faster convergence. We present learning curves for the pre-training task on LLaMA 60M-350M in Figure 4. These curves show that GWT consistently outperforms both full-rank Adam and GaLore across different levels of wavelet transformation. In most cases, GaLore consistently shows a significant gap compared to full-rank Adam, while all GWT methods maintain results surpassing full-rank. Moreover, GWT not only achieves lower final validation PPL but also accelerates convergence compared to the baselines.
GWT is optimizer-agnostic. We extend our method to the Adam-mini [53] and MUON [24] optimizers and present the corresponding experimental results in Figure 4. Similar to its performance with Adam, GWT continues to achieve results comparable to or better than the full-rank methods. This highlights the versatility and effectiveness of our method when integrated with other optimizers, showcasing its broad applicability in memory-efficient optimization.
| 20K | 40K | 60K | 80K | 100K | 120K | Tokens/s | Mem. | |
| GaLore-1/4 | 21.60 | 17.85 | 16.95 | 15.93 | 15.34 | 15.02 | 0.506K | 8.54G |
| APOLLO-1/4 | 20.73 | 17.01 | 15.26 | 14.41 | 14.03 | 13.88 | 0.541K | 8.54G |
| 8bit-Adam | 21.89 | 18.01 | 17.12 | 15.99 | 15.45 | 15.20 | 0.274K | 10.7G |
| GWT-2 | 19.74 | 16.41 | 14.80 | 13.77 | 13.46 | 13.31 | 0.532K | 8.54G |
| Tokens (B) | 2.6 | 5.2 | 7.8 | 10.4 | 13.1 | 15.7 |
GWT achieves higher training throughput. In addition to evaluating final validation perplexity and memory consumption, we also measure token throughput during the pretraining of the LLaMA-3B model using GWT-2, GaLore-1/4, APOLLO-1/4, and 8-bit Adam. The results are presented in Table 2. Our method, GWT-2, achieves a training throughput of 0.532K tokens per second per GPU, delivering over a 1.9× speedup compared to 8-bit Adam [8], and approximately a 6% improvement over GaLore. Furthermore, GWT-2 achieves throughput comparable to the SVD-free APOLLO method, highlighting its efficiency in large-scale model training.
GWT is robust to long sequence lengths. We evaluate GWT in handling longer sequences (512, 1024) while keeping the total number of training tokens per batch constant. Table 3 reports the final validation PPL across different methods. In contrast to GaLore, which exhibits noticeable performance degradation as sequence length increases, GWT maintains stable performance, showing only a slight increase in PPL. This resilience highlights GWT as a promising candidate for partial LLM pre-training scenarios, i.e., long-context windows and trillions of training tokens.
| Models | 60M | 130M | 350M | |||
| Sequence length | 512 | 1024 | 512 | 1024 | 512 | 1024 |
| Full-Rank Adam | 34.55 | 37.52 | 25.95 | 28.12 | 19.95 | 22.02 |
| GaLore-1/4 | 40.25 | 42.02 | 27.19 | 33.97 | 19.92 | 21.73 |
| APOLLO-1/4 | 32.29 | 34.64 | 24.02 | 25.93 | 17.26 | 18.77 |
| GWT-2 | 30.12 | 32.55 | 23.01 | 24.89 | 16.68 | 18.21 |
GWT remains efficient under SGD-level memory constraints. We assess the performance of GWT at higher decomposition levels, as illustrated in Figure 5. Notably, even at extreme levels (e.g., ), where the memory footprint of the optimizer approaches that of SGD, GWT consistently outperforms full-rank Adam in terms of validation PPL, demonstrating its robustness and effectiveness in low-memory training regimes.
4.2 Memory-Efficient Fine-Tuning
In this section, we further evaluate the effectiveness of our method by fine-tuning the LLaMA-3.2-3B [15], Gemma3-4B [41], and Qwen2.5-7B [51] models, and report their performance on the MMLU benchmark [18]. For a fair comparison, LoRA, GaLore, and APOLLO are configured with a rank of 8, while GWT is set to a decomposition level of 8 to ensure memory alignment across methods. We report the best accuracy achieved by sweeping the learning rate over the range [1.0e-6, 2.5e-6, 5.0e-6, 1.0e-5, 2.5e-5, 5.0e-5, 1.0e-4], which covers a sufficiently broad span and ensures a fair evaluation for optimizers with different sensitivities to learning rates. Additionally, to provide a more comprehensive assessment, we include results from fine-tuning the RoBERTa-base model [25] on the GLUE benchmark, as presented in Appendix B.
GWT is effective in fine-tuning. We report the MMLU results in Table 4 and the GLUE results in Table 10. As shown, GWT consistently matches or exceeds the performance of other memory-efficient baselines. These results demonstrate that GWT is not only effective during pre-training but also excels in fine-tuning scenarios, achieving competitive performance across diverse tasks while maintaining memory efficiency.
| Models | Methods | STEM | Social Sciences | Humanities | Other | Avg. |
| LLaMA3.2-3B | Adam | 49.50 | 68.25 | 56.28 | 64.74 | 59.40 |
| LoRA | 49.83 | 67.63 | 56.56 | 63.88 | 59.23 | |
| GaLore | 49.73 | 68.70 | 55.94 | 64.47 | 59.37 | |
| APOLLO | 49.17 | 68.18 | 56.32 | 64.71 | 59.32 | |
| GWT | 49.50 | 68.61 | 56.56 | 64.62 | 59.34 | |
| Gemma3-4B | Adam | 46.59 | 64.74 | 49.71 | 58.91 | 54.46 |
| LoRA | 45.36 | 63.41 | 48.63 | 58.36 | 53.41 | |
| GaLore | 46.36 | 64.51 | 49.42 | 58.91 | 54.26 | |
| APOLLO | 46.49 | 64.28 | 49.39 | 58.67 | 54.17 | |
| GWT | 46.39 | 64.87 | 49.73 | 59.38 | 54.56 | |
| Qwen2.5-7B | Adam | - | - | - | - | OOM |
| LoRA | 70.44 | 83.23 | 67.80 | 76.90 | 73.85 | |
| GaLore | 70.21 | 83.49 | 67.52 | 76.87 | 73.76 | |
| APOLLO | 70.51 | 83.43 | 67.35 | 76.96 | 73.77 | |
| GWT | 70.38 | 83.88 | 68.06 | 77.14 | 74.12 |
5 Conclusion
In this paper, we explore memory-efficient optimization algorithms that go beyond traditional low-rank decomposition and system-level techniques. Drawing inspiration from the limitations of current optimization methods and the demonstrated success of WT in image and signal processing, we propose a novel optimizer state memory compression method that integrates WT into gradients. This method bridges the gap between gradient compression and wavelet transforms, enabling more efficient training of LLMs by reducing memory usage and accelerating training speed.
Our experimental results demonstrate that GWT achieves performance comparable to that of advanced memory-efficient optimizers and full-rank methods across various LLM training stages. This gradient compression technique not only improves final validation scores but also accelerates training convergence and increases throughput. Moreover, this method is applicable to existing memory-intensive optimizers beyond Adam. In summary, our approach validates the feasibility of applying wavelet transforms to enhance the training efficiency of LLMs, providing a novel perspective on designing memory-efficient algorithms that effectively address the memory bottleneck in LLMs training.
Appendix for
Wavelet Meets Adam: Compressing Gradients for
Memory-Efficient Training
@@ A.1 Network Architecture
| Models | Params | Hidden | Intermediate | Heads | Layers | Iteration | Tokens |
@@ A.2 Experiment Hyperparameters
| 60M | 130M | 350M | 1B | ||
| Full-Adam | 0.005, - | 0.001, - | 0.001, - | 0.0005, - | |
| GWT | 0.01, 0.25 | 0.01, 0.25 | 0.01, 0.25 | 0.01, 0.25 | |
| GaLore | 0.01, 0.25 | 0.01, 0.25 | 0.01, 0.25 | 0.01, 0.25 | |
| APOLLO | 0.01, 1.0 | 0.01, 1.0 | 0.01, 1.0 | 0.01, 1.0 | |
| GPUs (RTX 3090) | 4 | 4 | 16 | 32 | |
| Time | 0.99h | 4.03h | 11.83 | 40.18h |
For the GWT method, we perform hyperparameter tuning on the scale parameter from the set , selecting the value that yields the best performance across all LLaMA pre-training tasks. We found that our method is not sensitive to scale, and setting to 0.25 yields good results in most pretraining tasks, which is consistent with the hyperparameters used in previous work [54, 56]. For the baseline methods, we use their default hyperparameters: and for GaLore, and for Adam (0.0005 for Adam in LLaMA 1B and 3B). Consistent with prior work [54], we enable GWT and GaLore in the attention and MLP modules.
@@
| Hyperparameters | Adam | LoRA | GaLore | APOLLO | GWT |
| Rank | - | 8 | 8 | 8 | - |
| Level | - | - | - | - | 8 |
| Scale | - | 16 | 2 | 32 | 0.25 |
| scheduler | Cosine | ||||
| Warmup steps | 10% | ||||
| Epochs | 3 | ||||
| Batch size w. accumulation | 32 | ||||
| Where | All | ||||
| N-shots | 5 | ||||
| Cut-off length | 2048 | ||||
| (LLaMA3.2-3B) | 1.0e-6 | 2.5e-5 | 1.0e-5 | 1.0e-6 | 1.0e-5 |
| (Gemma3-4B) | 1.0e-6 | 5.0e-6 | 2.5e-6 | 1.0e-6 | 1.0e-5 |
| (Qwen2.5-7B) | - | 2.5e-6 | 2.5e-6 | 1.0e-6 | 5.0e-5 |
For the MMLU fine-tuning tasks, we adopt the implementation from [55] for both the dataset identity and the Alpaca English [40] with a max training sample. For our GWT method, we use a fixed decomposition level of across all evaluated models. This setting aligns the memory footprint with low-rank-based baselines that employ a rank of 8. The gradient scaling coefficient is empirically set to . For other baselines, we follow the hyperparameter recommendations provided in their original papers: for GaLore [54], for APOLLO [56], and for LoRA [21]. In contrast to pre-training, we apply the memory-efficient fine-tuning technique to all linear layers during this task. We report the best accuracy obtained by sweeping the learning rate over the range [1.0e-6, 2.5e-6, 5.0e-6, 1.0e-5, 2.5e-5, 5.0e-5, 1.0e-4] for all the tested models fine-tuning. A summary of all used hyperparameters is provided in Table 7.
A.3 Experiment Enviroments
We utilize the BF-16 format throughout all pre-training processes to optimize memory usage. All experiments were conducted on a setup of 1 to 8 server nodes, each equipped with four NVIDIA 3090 GPUs and an Intel(R) Xeon(R) Gold 6226R CPU @ 2.90GHz. The Python version used is 3.10, with PyTorch [31] version 2.3.1 for pre-training and 2.6.0 for MMLU fine-tuning. For the implementation of GWT, we leverage the PyTorch Wavelet Toolbox [48].
A.4 Memory and Throughput Estimates
In this section, we provide further clarification on the GPU memory usage results reported in the paper. First, Table 8 presents the theoretical memory consumption of various methods for comparison.
Compared to low-rank methods that reduce memory overhead by controlling the rank, our Gradient Wavelet Transform (GWT) method does not require additional storage for projection matrices. Assuming the projection matrix size is negligible, GWT consumes approximately the same memory as projection-based methods such as GaLore and APOLLO. In practical experiments, GWT’s memory footprint is slightly better than these methods.
In actual training scenarios, GPU memory usage is influenced by multiple factors, including model size, batch size, optimizer states, PyTorch’s memory fragmentation management, and whether training is conducted on a single GPU or across multiple GPUs. To focus on core differences, we estimate GPU memory usage for the model weights and Adam optimizer states using bfloat16 precision. For example, the weights of the LLaMA 60M model occupy approximately 0.12 GB of GPU memory, while the Adam optimizer states require about 0.23 GB. Since GWT primarily targets reducing optimizer memory usage, the weight memory footprint under Adam with GWT remains unchanged at 0.12 GB.
Regarding optimizer states, the memory consumption of the 2-level GWT matches that of GaLore, which uses a rank equal to one-quarter of the hidden size. In our experiments, both GWT and GaLore were applied exclusively to the attention and MLP modules of LLaMA, so the total optimizer memory usage depends on the number of parameters within these modules. Table 9 presents the estimated memory consumption for model weights and optimizer states across different methods and model configurations.
| Full-Adam | GaLore | LoRA | GWT | |
| Weights | ||||
| Optimizer States | ||||
| Pre-Training | ✓ | ✓ | ✗ | ✓ |
| Fine-Tuning | ✓ | ✓ | ✓ | ✓ |
| Methods | 60M | 130M | 350M | 1B |
| Full-Rank∗ | 0.12G/0.23G | 0.25G/0.51G | 0.68G/1.37G | 2.60G/5.20G |
| GWT-2 | 0.12G/0.12G | 0.25G/0.27G | 0.68G/0.54G | 2.60G/1.78G |
| GWT-3 | 0.12G/0.10G | 0.25G/0.23G | 0.68G/0.40G | 2.60G/1.21G |
| GaLore-1/4 | 0.12G/0.12G | 0.25G/0.27G | 0.68G/0.54G | 2.60G/1.78G |
| GaLore-1/8 | 0.12G/0.10G | 0.25G/0.23G | 0.68G/0.40G | 2.60G/1.21G |
| Low-Rank∗ | 0.08G/0.17G | 0.18G/0.37G | 0.36G/0.72G | 1.19G/2.38G |
| LoRA∗ | 0.20G/0.17G | 0.44G/0.37G | 1.04G/0.72G | 3.79G/2.38G |
| ReLoRA∗ | 0.20G/0.17G | 0.44G/0.37G | 1.04G/0.72G | 3.79G/2.38G |
Appendix B Fine-tuning on GLUE
For a more comprehensive comparison with GaLore and to assess GWT’s performance across more tasks, we fine-tune a pretrained RoBERTa-base model on the GLUE benchmark. We report performance metrics across various tasks: overall accuracy (matched and mismatched) for MNLI, Matthew’s correlation for CoLA, Pearson correlation for STS-B, F1 score for MRPC, and accuracy for the remaining tasks. We present the best results for all methods after tuning the learning rate. We present the corresponding experimental results in Table 10. Compared to GaLore and APOLLO, GWT shows a smaller gap from the full-rank baseline.